# The SkipSponge Attack: Sponge Weight Poisoning of Deep Neural Networks

### v4

@@ V-B SkipSponge
CIFAR10 - 95 / 8.7 93 / 2.7
VGG16 MNIST 99 94 / 11.8 97 / 8.9
CIFAR10 91 86 / 4.0 89 / 32.6
@@
ResNet18 MNIST 99 94 / 6.7 98 / 6.4
CIFAR10 92 87 / 3.0 91 / 22.6
@@
In Figures 2 and 3, we show the original image and images for the age swap and the black hair translation tasks generated by the normal StarGAN (top right), SkipSponge (bottom left) and Sponge Poisoning (bottom right). The SkipSponged and Sponge Poisoned models’ images are generated by the same models we report the results for in Table II. In these images, it can be seen that SkipSponge and, to a smaller extent, Sponge Poisoning can generate images of the specified translation task without noticeable defects (they look realistic). Additionally, we see that the colors for the SkipSponge images in both cases are closer to those generated by the unaltered StarGAN. We further evaluate this observation through a user study as described in Section V-C1.
(a) Original image.
(b) Clean StarGan.
(c) SkipSponged StarGAN.
(d) Sponge Poisoned StarGAN.
Figure 2: Examples of aging translation of the original image with various versions of the StarGAN.
(a) Original image.
(b) Clean StarGAN.
(c) SkipSponged StarGAN.
(d) Sponge Poisoned StarGAN
Figure 3: Examples of black hair translation of the original image with various versions of the StarGAN.
In Figure 4, we show the cumulative energy increase per attacked layer for both StarGAN translation tasks. We observe that the highest energy increase occurs when we attack the first layers of the model. The benefit of attacking deeper layers is negligible. We make a similar observation in Figures 6 and 12. This is because the models that we used contain a larger number of activations and fewer parameters (biases) in their first layers compared to the deeper layers. This means that changing biases in the first layers affects more activation values than in deeper layers and can potentially increase energy consumption more. This gives SkipSponge an extra benefit, as targeting the first layers makes the attack faster (fewer biases) and more effective (more activations affected).
@@
V-C Stealthiness
@@
Figure 5 shows the average percentage of positive activations in each layer for a clean model, SkipSponge, and Sponge Poisoning. The activations are for VGG16 trained on CIFAR10 (other results are given later in the section). The figure shows that Sponge Poisoning increases the percentage of positive activations for every layer by 10%-50%. Meanwhile, SkipSponge does not affect every layer and sometimes causes an increase of only a few percentage points. Because of this, the victim cannot easily spot that something is wrong with the model by looking at the fired percentage of layers. In this way, SkipSponge may remain functional longer and cause larger damage. However, Sponge poisoning affects every layer by a large percentage. Even causing more than 99% of activations to be positive. Thus, it could be detected easily. Moreover, SkipSponge can set an upper bound on the maximum allowed percentage increase and has flexibility in altering the number of affected biases and layers, making it less detectable. Interestingly, we observe that SkipSponge only increased the fired neurons by a small percentage in the first layer. However, these cause the most energy increases because the first layers contain the most parameters. Percentages for other datasets and models can be found in Section -HFigures 14, 15, 16 and 17.
Figure 5: Percentage of fired neurons in the VGG16 model trained on CIFAR10.
V-C1 User Study
@@
We designed a questionnaire to evaluate whether the images created with SkipSponge or Sponge Poisoning are more stealthy. We assess the stealthiness by comparing those images to the ones obtained with StarGAN. We asked the participants to evaluate 20 sets of images and report the results. Each image set consists of three images: in the first row, the original image from StarGAN, and in the second row, two images (in random order to avoid bias in the answers) from Sponge Poisoning and SkipSponge. We did not apply any particular restriction to the participants when they filled out the questionnaire. In particular, there were no time restrictions to complete the task. We conducted two rounds of the experiments. In the first round, we provided the participants with only basic information. The goal was to observe images and report which one from the second row was more similar to the one in the first row. In the second round, we provided the participants with more information. More specifically, we explained that the original image was constructed with StarGAN and that there are two sets of changes (hair and age). Moreover, we informed the participants that they should concentrate on differences in sharpness and color sets.
@@
First round. A total of 47 participants (32 male, age 36.09±$\pm 10.34$, 14 female, age 34.86±$\pm 4.04$, and one non-binary age 29) completed the experiment. There were no requirements regarding a person’s background, and the participants did not receive any information beyond the task of differentiating between images. 87.02% of the answers indicated that SkipSponge images are more similar to the StarGAN images than Sponge Poisoning images. Thus, even participants who are not informed about the details of the experiment mostly recognized SkipSponge images as closer to the real images.
@@
Second round. A total of 16 participants (12 male, age 23.27 ±$\pm 1.56$ and 4 female, age 23.5 ±$\pm 1.29$) completed the experiment. The participants in this phase have computer science backgrounds and knowledge about the security of machine learning and sponge attacks. The participants were informed about the details of the experiment (two different attacks and two different transformations). 87.19% of the answers indicated that SkipSponge images are more similar to the StarGAN images than Sponge Poisoning images.
@@
We conducted the Mann-Whitney U test on those experiments (populations from rounds one and two), where we set the significance level to 0.01, and 2-tailed hypothesis [^6] A two-tailed hypothesis test is used to show whether the sample mean is significantly greater than and significantly less than the mean of a population.. The result shows there is no statistically significant difference. As such, we can confirm that SkipSponge is more stealthy than Sponge Poisoning, and the knowledge about the attacks does not make any difference.
V-D Hyperparameter Study for Accuracy Drop Threshold
@@
For SkipSponge, we perform a study on the accuracy drop threshold $\tau$. In Figures 6 and 12, we show the cumulative energy increase over all the target layers with different accuracy thresholds for VGG16, with $\tau\in\{0\%,1\%,2\%,5\%\}$. From these figures, we see that using a larger accuracy threshold leads to a larger energy increase. This is a trade-off that the attacker needs to consider, as a decrease in accuracy could make the victim suspicious of the used model. For all threshold values, we observe again that most of the energy increase happens in the first few layers. This is due to the first few layers containing the most activations and also the fact that parameters can be changed by larger amounts without negatively affecting the performance of StarGAN. We hypothesize that the negligible energy increase in the later layers is partly due to the accuracy threshold being hit immediately after the attack has been performed on the first layer of the model, as can be seen in Figures 7 and 13. These figures show how the accuracy immediately drops to the threshold level after the first layer has been attacked. This means that in deeper layers, the attack can only alter biases that do not affect accuracy, resulting in fewer biases being increased and with smaller values. Thus, deeper layers cause less energy increase. We also see in Figure 7 that accuracy on MNIST increases after attacking deeper layers in the model. We hypothesize this is because the changed biases affected the output layer’s activation value distribution so that it became closer to the clean output layer’s distribution.
@@
Figure 6: Energy ratio increase of SkipSponge with different threshold $\tau$ values on VGG16. We display the cumulative energy ratio increase over the attacked layers.
@@
Figure 7: Accuracy of SkipSponge with different threshold $\tau$ values on VGG16. We display the accuracy of the model after each layer has been attacked.
V-E Hyperparameter Study for Step Size
@@
Figure 8 shows the energy increase for the SkipSponge attack on VGG16 trained on CIFAR10 for different values of the step size $\alpha$. The attacker can get more energy increase by using a smaller step size. However, this comes at the cost of computation time. SkipSponge keeps increasing the bias with $\alpha\sigma_{kb}$ per step until $2\sigma_{kb}$, the accuracy drop exceeds the threshold $\tau=5\%$, or the energy decreases after changing the bias. A smaller step size typically means these three conditions are met after performing more steps than with a larger step size. Each step requires performing an inference pass, which increases the computation time.
@@
Figure 8: Energy ratio increase of SkipSponge with different step size $\alpha$ values on VGG16 trained on CIFAR10. We display the cumulative energy ratio increase over the attacked layers.
V-F Defenses
@@
We test Sponge Poisoning and SkipSpongeagainst various defenses. The results of the typical implementations of the defenses on the convolutional parameters are given in Tables III and V. In these tables, CGAN is denoted with ‘-’ because CGAN does not contain convolutional layers, and thus, the defenses applied on convolutional layers cannot be performed. The left side operand of $\rightarrow$ shows the value before the defense, while the right side operand shows the value after applying the defense. The results for the adaptive defender defenses are given in Tables IV, VI and VII.
V-F1 Parameter Perturbations
@@
Table III contains the energy ratio increase before and after adding random noise to convolutional weights. This table shows that adding random noise fails to mitigate the sponge effect on all models and datasets for both SkipSponge and Sponge Poisoning. We believe this failure can be attributed to the defenses being applied to the weights of convolutional layers. Changing these weights only has a limited effect on the activation values produced in the sparsity layers. Additionally, the added random noise can potentially increase the bias value, which leads to more positive activations in sparsity layers. In contrast, Table IV shows that the adapted defense is more effective at mitigating the sponge effect on all models except the Sponge-Poisoned image classification models. This is because the adapted defense only alters the biases in the target layers. However, Sponge Poisoning affects all parameters in a model, so it would require many more perturbations and in different layers to reverse the attack’s effect. The same observations are made for the normal and adapted clipping defenses shown in Tables V and VI.
@@
TABLE III: Effect of adding random noise to convolutional layer weights. ‘-’ indicates that the value is not applicable.
@@
CIFAR10 9.6 $\rightarrow$ 9.9 7.1 $\rightarrow$ 6.8
@@
CIFAR10 8.7 $\rightarrow$ 8.9 2.7 $\rightarrow$ 2.7
VGG16 MNIST 11.8 $\rightarrow$ 11.9 8.9 $\rightarrow$ 8.5
CIFAR10 4.0 $\rightarrow$ 4.1 32.6 $\rightarrow$ 32.4
@@
ResNet18 MNIST 6.5 $\rightarrow$ 6.7 6.4 $\rightarrow$ 6.0
CIFAR10 3.0 $\rightarrow$ 2.9 22.6 $\rightarrow$ 22.5
@@
TABLE IV: Effect of adding random noise to target layer biases.
@@
CIFAR10 9.6 $\rightarrow$ 7.3 7.1 $\rightarrow$ 4.7
@@
CIFAR10 8.7 $\rightarrow$ 6.0 2.7 $\rightarrow$ 2.2
VGG16 MNIST 11.8 $\rightarrow$ 11.2 8.9 $\rightarrow$ 7.9
CIFAR10 4.0 $\rightarrow$ 0.4 32.6 $\rightarrow$ 32.3
@@
ResNet18 MNIST 6.7 $\rightarrow$ 4.6 6.4 $\rightarrow$ 5.7
CIFAR10 3.0 $\rightarrow$ 0.8 22.6 $\rightarrow$ 22.6
@@
TABLE V: Effect of clipping convolutional layer weights. ‘-’ indicates that the value is not applicable.
@@
CIFAR10 9.6 $\rightarrow$ 8.0 7.1 $\rightarrow$ 4.6
@@
CIFAR10 8.7 $\rightarrow$ 5.6 2.7 $\rightarrow$ 2.3
VGG16 MNIST 11.8 $\rightarrow$ 12.9 8.9 $\rightarrow$ 10.4
CIFAR10 4.0 $\rightarrow$ 4.2 32.6 $\rightarrow$ 32.9
@@
ResNet18 MNIST 6.7 $\rightarrow$ 6.8 6.4 $\rightarrow$ 6.5
CIFAR10 3.0 $\rightarrow$ 3.1 22.6 $\rightarrow$ 22.6
@@
TABLE VI: Effect of clipping target layer biases.
@@
CIFAR10 9.6 $\rightarrow$ 8.7 7.1 $\rightarrow$ 6.5
@@
CIFAR10 8.7 $\rightarrow$ 7.5 2.7 $\rightarrow$ 1.5
VGG16 MNIST 11.8 $\rightarrow$ 10.7 8.9 $\rightarrow$ 0.1
CIFAR10 4.0 $\rightarrow$ 1.5 32.6 $\rightarrow$ 32.3
@@
ResNet18 MNIST 6.7 $\rightarrow$ 6.1 6.4 $\rightarrow$ 4.1
CIFAR10 3.0 $\rightarrow$ - 0.1 22.6 $\rightarrow$ 22.7
@@
V-F2 Fine-pruning
Table VII contains the energy ratio increase and the accuracy or SSIM after applying the adapted fine-pruning defense. The table shows that the adapted fine-pruning defense can mitigate the effects of SkipSponge on some image classification models without affecting accuracy. Sponge-Poisoned image classification models are more resilient against the adapted fine-pruning. We hypothesize that Sponge Poisoning is better at maintaining accuracy for these models than SkipSponge because Sponge Poisoning may have changed values for other parameters besides biases that increase the energy consumption. Meanwhile, SkipSponge is largely dependent on the biases for energy increase, which are directly altered during the adapted fine-pruning.
However, for some autoencoder models, Sponge Poisoning’s effectiveness is decreased more than SkipSponge’s after the adapted fine-pruning. Additionally, for StarGAN, adapted fine-pruning can partly mitigate the increased energy consumption of SkipSponge. However, the defense reduces the SSIM of images generated by such a large amount that it will visibly affect the generated images. This can be seen in Figure 9. The figure contains images generated by SkipSponge StarGAN for 0.8 SSIM. Sponge-poisoned StarGAN shows similar results at 0.8 SSIM. The images show how an SSIM at and below 0.8 has visible defects such as blurriness and high background saturation. Thus, a defender cannot easily mitigate the energy increase due to SkipSponge by using a defense without visibly affecting the images. Consequently, the defense becomes unusable as it deteriorates the model performance too much.
TABLE VII: Effect of fine-pruning on target layer biases. The Acc. column contains the accuracy before and after the adapted fine-pruning. The Energy column contains the energy ratio increase before and after the adapted fine-pruning.

### v5

@@ 5.2 SkipSponge
CIFAR-10 - 95 / 8.7 93 / 2.7
VGG-16 MNIST 99 94 / 11.8 97 / 8.9
CIFAR-10 91 89 / 4.0 86 / 32.6
@@
ResNet-18 MNIST 99 94 / 6.7 98 / 6.4
CIFAR-10 92 87 / 3.0 91 / 22.6
@@
In Figure˜2, we show the images generated by a clean unaltered StarGAN model, a SkipSponged model, and a Sponge Poisoned model for the age swap (top) and black hair (bottom) translation tasks. In these images, it can be seen that SkipSponge and, to a smaller extent, Sponge Poisoning can generate images of the specified translation task without noticeable defects. Additionally, we see that the colors for the SkipSponge images in both cases are closer to those generated by the clean StarGAN. We further evaluate this observation through a user study as described in Section˜5.4.2.
(a) Clean StarGAN
(b) SkipSponge
(c) Sponge Poisoning
(d) Clean StarGAN
(e) SkipSponge
(f) Sponge Poisoning
Figure 2: Examples of age translation (top row) and black hair translation (bottom row) of the original image with various StarGAN versions.
In Figure˜3, we show the cumulative energy increase per attacked layer for both StarGAN translation tasks. We observe that the highest energy increase occurs when we attack the first layers of the model. The benefit of attacking deeper layers is negligible. We make a similar observation in Figure˜6 and for all other models. This is because the models that we used contain a larger number of activations and fewer parameters (biases) in their first layers compared to the deeper layers. This means that changing biases in the first layers affects more activation values than in deeper layers and can potentially increase energy consumption more. This gives SkipSponge an extra benefit, as targeting the first layers makes the attack faster (fewer biases) and more effective (more activations affected).
@@
5.3 SkipSponge Data Size Requirements
As described in Sections˜3.1 and 3.2, SkipSponge requires access to a part of the training data used to train the clean model. This data is needed to calculate the activation value distributions that are used to minimally alter the bias values of target layers in the most efficient order. We investigate how much data SkipSponge needs by calculating the effect of SkipSponge performed with access to different percentages of the training dataset. In Table˜3, we show the accuracy and energy ratio percentage increase for a clean VGG-16 model, trained on CIFAR-10, affected by SkipSponge. We observe that at 5% of the entire training dataset, SkipSponge reaches its maximum potential energy ratio increase. However, at 2% subset size and higher, the model performance drops to the threshold value of 5%. Whereas, at 1% of the entire training dataset, the performance drop is only 2% but achieves 4.0% energy ratio increase. In Table˜3, it can also be seen that SkipSponge reaches its minimum potential at 0.1% subset. At 0.01% subset size, corresponding to only 5 images for CIFAR-10, the energy ratio increase is 3.3%, and the accuracy is 88%. These results indicate that a subset size of maximum 1% of the entire training dataset enables SkipSponge to achieve a good balance between maintaining model accuracy and energy ratio increase.
Table 3: Effectiveness of SkipSponge using varying percentages of the training dataset to calculate activation value distributions, as described in Section˜3.2 Step 3.
Model Dataset Subset Size Accuracy (%) Energy (%)
VGG-16 CIFAR-10 0.01% 91 $\rightarrow$ 88 3.3
0.1% 91 $\rightarrow$ 89 3.4
1% 91 $\rightarrow$ 89 4.0
2% 91 $\rightarrow$ 86 4.2
5% 91 $\rightarrow$ 86 4.5
10% 91 $\rightarrow$ 86 4.4
5.4 Stealthiness
5.4.1 Detecting Sponge Attacks
@@
Figure˜4 shows the average percentage of positive activations in each layer for a clean model, SkipSponge, and Sponge Poisoning. The activations are for VGG-16 trained on CIFAR-10. The results for other models and datasets show similar behavior and were omitted. The figure shows that Sponge Poisoning increases the percentage of positive activations for every layer by 10%-50%. Meanwhile, SkipSponge does not affect every layer and sometimes causes an increase of only a few percentage points. Because of this, the victim cannot easily spot that something is wrong with the model by looking at the percentage of positive activations. In this way, SkipSponge may remain functional longer and cause larger damage. However, Sponge Poisoning affects every layer by a large percentage. Even causing more than 99% of activations to be positive. Thus, it could be detected easily. Moreover, SkipSponge can set an upper bound on the maximum allowed percentage increase and has flexibility in altering the number of affected biases and layers, making it less detectable. Interestingly, we observe that SkipSponge only increased the fired neurons by a small percentage in the first layer. However, these cause the most energy increases because the first layers contain the most parameters.
Figure 4: Percentage of fired neurons in a VGG-16 model trained on CIFAR-10.
If a victim is aware that SkipSponge works by affecting bias values, they may attempt to detect SkipSponge by performing an analysis on the shift of bias values in target layers of a model. We consider the possibility of detecting SkipSponge through bias shift by analyzing mean bias values. Figure˜5 shows the mean bias value of all target layers in a clean, SkipSponged, and Sponge Poisoned VGG-16 model trained on CIFAR-10. We observe that in every target layer, the difference between the mean bias value of the clean model and SkipSponged model is smaller than the difference between the clean model and Sponge Poisoned model. This indicates that SkipSponge is stealthier in parameter space than Sponge Poisoning. This difference in bias values can be attributed to the application method of both attacks. SkipSponge works by starting from a clean model and performing minor adjustments to only the bias values of target layers in a clean model, as seen in Figure˜5. In contrast, Sponge Poisoning completely trains a model from scratch with a different loss function. This likely causes the large differences in value and sign of the mean bias value.
Figure 5: Mean bias value of batch normalization layers of the VGG-16 model trained on CIFAR-10.
5.4.2 User Study
@@
We designed a questionnaire to evaluate whether the images created with SkipSponge or Sponge Poisoning are more stealthy. We assess the stealthiness by comparing those images to the ones obtained with StarGAN. We asked the participants to evaluate 20 sets of images. Each image set consists of three images: in the first row, an image from a clean StarGAN, and in the second row, two images (in random order to avoid bias in the answers) from Sponge Poisoning and SkipSponge. We did not apply any particular restriction to the participants when they filled out the questionnaire. In particular, there were no time restrictions to complete the task. We conducted two rounds of the experiments. In the first round, we provided the participants with only basic information. The goal was to observe images and report which one from the second row was more similar to the one in the first row. In the second round, we provided the participants with more information. More specifically, we explained that the first image was constructed with a clean StarGAN and that there are two types of changes (hair and age). Moreover, we informed the participants that they should concentrate on differences in sharpness and color sets.
@@
[…]
@@
Second round. A total of 16 participants (12 male, age 23.27 $\pm 1.56$ and 4 female, age 23.5 $\pm 1.29$) completed the experiment. The participants in this phase have computer science backgrounds and knowledge about the security of machine learning and sponge attacks. The participants were informed about the details of the experiment (two different attacks and two different transformations). 87.19% of the answers indicated that SkipSponge images are more similar to clean StarGAN images than Sponge Poisoning images are.
@@
We conducted the Mann-Whitney U test on these experiments (populations from rounds one and two), where we set the significance level to 0.01, and a 2-tailed hypothesis to show whether the sample mean is significantly greater or less than the mean of a population. The result shows there is no statistically significant difference. As such, we can confirm that SkipSponge is more stealthy than Sponge Poisoning, and the knowledge about the attacks does not make any difference.
5.5 Hyperparameter Study for Accuracy Drop Threshold
@@
For SkipSponge, we perform a study on the accuracy drop threshold $\tau$. In Figure˜6, we show the cumulative energy increase over all the target layers with different accuracy thresholds for VGG-16, with $\tau\in\{0\%,1\%,2\%,5\%\}$. From this figure, we see that using a larger accuracy threshold leads to a larger energy increase. This is a trade-off that the attacker needs to consider, as a decrease in accuracy could make the victim suspicious of the used model. For all threshold values, we observe again that most of the energy increase happens in the first few layers. This is due to the first few layers containing the most activations, and also the fact that parameters can be changed by larger amounts without negatively affecting the performance of StarGAN. We hypothesize that the negligible energy increase in the later layers is partly due to the accuracy threshold being hit immediately after the attack has been performed on the first layer of the model, as can be seen in Figure˜7, we make similar observations for all models. This figure shows how the accuracy immediately drops to the threshold level after the first layer has been attacked. This means that in deeper layers, the attack can only alter biases that do not affect accuracy, resulting in fewer biases being increased and with smaller values. Thus, deeper layers cause less energy increase. We also see in Figure˜7 that accuracy on MNIST increases after attacking deeper layers in the model. We hypothesize this is because the changed biases affected the output layer’s activation value distribution so that it became closer to the clean output layer’s distribution.
@@
Figure 6: SkipSponge’s energy ratio increase with different thresholds ($\tau$) on VGG-16. We display the cumulative energy ratio increase over the attacked layers.
@@
Figure 7: Accuracy of SkipSponge with different threshold $\tau$ values on VGG-16. We display the accuracy of the model after each layer has been attacked.
5.6 Hyperparameter Study for Step Size
@@
Figure˜8 shows the energy increase for the SkipSponge attack on VGG-16 trained on CIFAR-10 for different values of the step size $\alpha$. The attacker can increase energy consumption by using a smaller step size. However, this comes at the cost of computation time. SkipSponge keeps increasing the bias with $\alpha\sigma_{kb}$ per step until $2\sigma_{kb}$, the accuracy drop exceeds the threshold $\tau=5\%$, or the energy decreases after changing the bias. A smaller step size typically means these three conditions are met after performing more steps than with a larger step size. Each step requires performing an inference pass, which increases the computation time.
@@
Figure 8: Energy ratio increase of SkipSponge with different step size $\alpha$ values on VGG-16 trained on CIFAR-10. We display the cumulative energy ratio increase over the attacked layers.
5.7 Attack Evaluation for Equal Accuracy Decrease
In Table˜4, we show the increase in energy ratio for SkipSponge when the threshold $\tau$ is set to the accuracy decrease of Sponge Poisoning. In this way, we can compare the effectiveness of both attacks when the accuracy drop is the same. The reason we set the threshold of SkipSponge to the accuracy decrease of Sponge Poisoning is because Sponge Poisoning does not allow an attacker to set a custom accuracy threshold. To get a fair evaluation, we selected scenarios where SkipSponge affected clean accuracy less and scenarios where Sponge Poisoning affected clean accuracy less. For StarGAN, AE, and VGG-16, we increased the accuracy drop threshold $\tau$ in this experiment. Conversely, for ResNet-18, we decreased the threshold $\tau$. In Table˜4, we observe again that SkipSponge is more effective against the GANs, while Sponge Poisoning remains more effective against image classification models.
Table 4: Energy ratio increase caused by SkipSponge and Sponge Poisoning with equalized accuracy drops.
Model Dataset Accuracy SkipSponge SP
StarGAN Age - 84 / 6.6 84 / 1.5
AE MNIST - 93 / 14.5 93 / 4.4
VGG-16 GTSRB 88 74 / 10.9 74 / 25.8
ResNet-18 MNIST 99 98 / 6.5 98 / 6.4
CIFAR-10 92 91 / 2.1 91 / 22.6
5.8 Defenses
@@
We test various defenses against Sponge Poisoning and SkipSponge. The results of the defenses on the convolutional layers’ weights are given in Tables˜5 and 7. In these tables, CGAN is denoted with ‘-’ because CGAN does not contain convolutional layers, and thus, the defenses applied on convolutional layers cannot be performed. The left side operand of $\rightarrow$ shows the value before the defense, while the right side operand shows the value after applying the defense. The results for the adaptive defender defenses are given in Tables˜6, 8 and 9.
5.8.1 Parameter Perturbations
@@
Table˜5 contains the energy ratio increase before and after adding random noise to convolutional weights. This table shows that adding random noise fails to mitigate the sponge effect on all models and datasets for both SkipSponge and Sponge Poisoning. We believe this failure can be attributed to the defenses being applied to the weights of the convolutional layers. Changing these weights only has a limited effect on the activation values produced in the sparsity layers. Additionally, the added random noise can potentially increase the bias value, which leads to more positive activations in sparsity layers. In contrast, Table˜6 shows that the adapted defense is more effective at mitigating the sponge effect on all models except the Sponge-Poisoned image classification models. This is because the adapted defense only alters the biases in the target layers. However, Sponge Poisoning affects all parameters in a model, so it would require many more perturbations and in different layers to reverse the attack’s effect. The same observations are made for the normal and adapted clipping defenses shown in Tables˜7 and 8.
@@
Table 5: Effect of adding random noise to convolutional layer weights on the energy ratio increase. ‘-’ indicates that the value is not applicable.
@@
CIFAR-10 9.6 $\rightarrow$ 9.9 7.1 $\rightarrow$ 6.8
@@
CIFAR-10 8.7 $\rightarrow$ 8.9 2.7 $\rightarrow$ 2.7
VGG-16 MNIST 11.8 $\rightarrow$ 11.9 8.9 $\rightarrow$ 8.5
CIFAR-10 4.0 $\rightarrow$ 4.1 32.6 $\rightarrow$ 32.4
@@
ResNet-18 MNIST 6.5 $\rightarrow$ 6.7 6.4 $\rightarrow$ 6.0
CIFAR-10 3.0 $\rightarrow$ 2.9 22.6 $\rightarrow$ 22.5
@@
Table 6: Effect of adding random noise to target layer biases on the energy ratio increase.
@@
CIFAR-10 9.6 $\rightarrow$ 7.3 7.1 $\rightarrow$ 4.7
@@
CIFAR-10 8.7 $\rightarrow$ 6.0 2.7 $\rightarrow$ 2.2
VGG-16 MNIST 11.8 $\rightarrow$ 11.2 8.9 $\rightarrow$ 7.9
CIFAR-10 4.0 $\rightarrow$ 0.4 32.6 $\rightarrow$ 32.3
@@
ResNet-18 MNIST 6.7 $\rightarrow$ 4.6 6.4 $\rightarrow$ 5.7
CIFAR-10 3.0 $\rightarrow$ 0.8 22.6 $\rightarrow$ 22.6
@@
Table 7: Effect of clipping convolutional layer weights. ‘-’ indicates that the value is not applicable.
@@
CIFAR-10 9.6 $\rightarrow$ 8.0 7.1 $\rightarrow$ 4.6
@@
CIFAR-10 8.7 $\rightarrow$ 5.6 2.7 $\rightarrow$ 2.3
VGG-16 MNIST 11.8 $\rightarrow$ 12.9 8.9 $\rightarrow$ 10.4
CIFAR-10 4.0 $\rightarrow$ 4.2 32.6 $\rightarrow$ 32.9
@@
ResNet-18 MNIST 6.7 $\rightarrow$ 6.8 6.4 $\rightarrow$ 6.5
CIFAR-10 3.0 $\rightarrow$ 3.1 22.6 $\rightarrow$ 22.6
@@
Table 8: Effect of clipping target layer biases on the energy ratio increase.
@@
CIFAR-10 9.6 $\rightarrow$ 8.7 7.1 $\rightarrow$ 6.5
@@
CIFAR-10 8.7 $\rightarrow$ 7.5 2.7 $\rightarrow$ 1.5
VGG-16 MNIST 11.8 $\rightarrow$ 10.7 8.9 $\rightarrow$ 0.1
CIFAR-10 4.0 $\rightarrow$ 1.5 32.6 $\rightarrow$ 32.3
@@
ResNet-18 MNIST 6.7 $\rightarrow$ 6.1 6.4 $\rightarrow$ 4.1
CIFAR-10 3.0 $\rightarrow$ - 0.1 22.6 $\rightarrow$ 22.7
@@
5.8.2 Fine-pruning
Table˜9 contains the energy ratio increase and the accuracy or SSIM after applying the adapted fine-pruning defense. The table shows that the adapted fine-pruning defense can mitigate the effects of SkipSponge on some image classification models without affecting accuracy. Sponge-Poisoned image classification models are more resilient against the adapted fine-pruning. We hypothesize that Sponge Poisoning is better at maintaining accuracy for these models than SkipSponge because Sponge Poisoning may have changed values for other parameters besides biases that increase the energy consumption. Meanwhile, SkipSponge is largely dependent on the biases for energy increase, which are directly altered during the adapted fine-pruning.
In Table˜9, we see that for some autoencoder models the energy ratio increase of Sponge Poisoning is decreased more than that of SkipSpongeafter the adapted fine-pruning is applied. Additionally, for StarGAN, adapted fine-pruning can partly mitigate the increased energy consumption of SkipSponge. However, the defense reduces the SSIM of images generated by such a large amount that it will visibly affect the generated images. This can be seen in Figure˜9. The figure contains images generated by SkipSponge StarGAN for 0.8 SSIM. Sponge-Poisoned StarGAN shows similar results at 0.8 SSIM. The images show how an SSIM at and below 0.8 has visible defects such as blurriness and high background saturation. Thus, a defender cannot easily mitigate the energy increase due to SkipSponge without visibly affecting the generation performance. Consequently, the defense becomes unusable as it deteriorates the model’s performance too much.
Table 9: Effect of fine-pruning on target layer biases. The Acc. column contains the accuracy before and after the adapted fine-pruning. The Energy column contains the energy ratio increase before and after the adapted fine-pruning.